# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ 3 Graduate School of Science
The aging law and the slip law are two representative evolution laws of the rate- and state-dependent friction (RSF) law, based on canonical behaviors in three types of laboratory experiments: slide-hold-slide (SHS), velocity-step (VS), and steady-state (SS) tests. The aging law explains the SHS canon but contradicts the VS canon, and vice versa for the slip law. The later proposed composite law, which switches these two laws according to the slip rate $V$, explains both canons but contradicts the SS canon. The present study constructs evolution laws satisfying all three canons throughout the range of variables where experiments have confirmed the canons. By recompiling the three canons, we have derived constraints on the evolution law and find that the evolution rates in the strengthening phases of the SHS and VS canons are so different that complete reconciliation throughout the entire range of variables is mathematically impossible. However, for the limited range of variables probed by experiments so far, we have found that the SHS and VS canons can be reconciled without violating the SS canon by switching the evolution function according to $\Omega$, the ratio of the state $\theta$ to its steady-state value $\theta_{\rm SS}$ for the instantaneous slip rate. We could generally show that, as long as the state evolution rate $\dot{\theta}$ depends only on the instantaneous values of $V$ and $\theta$, simultaneous reproduction of the three canons, throughout the experimentally confirmed range, requires the aging-law-like evolution for $\Omega$ sufficiently below a threshold $\beta$ and the slip-law-like evolution for $\Omega$ sufficiently above $\beta$. The validity of the canons in existing experiments suggests $\beta\lesssim 0.01$.
@@ 1 Introduction
Formulas that can describe the consensual frictional behaviors in three representative types of experiments (eq. 26, eq. 32, and eqs. 48 and 49) have been pursued since Dieterich (1979). Dieterich’s formulation (1979) is primarily motivated by SS and SHS tests by Dieterich (1972) and Dieterich (1978). It was reorganized into a differential equation (eq. 4) by Ruina (1983), called the aging law (Dieterich law) today. However, the aging law contradicts VS tests. By linear expansion of the aging law around the steady states, Ruina (1983) proposed another differential equation (eq. 5) called the slip law (Ruina law) today. Ruina (1983) pointed out that the slip law is a straightforward translation of VS and SS behaviors into a differential equation. However, the slip law cannot describe the log-time healing observed under low shear stresses where slip is negligible, although it can explain another, more popular type of SHS (the stationary-loadpoint SHS, §3.3) where shear stress remains pretty high, allowing for significant slip to continue in its quasi-static hold period. Thus, neither law has succeeded in compiling the different types of experiments. This long-standing discrepancy, the aging vs. slip problem (Beeler et al., 1994; Ampuero & Rubin, 2008), is the background of the present paper. Before starting the lengthy mathematical derivation, we summarize the requirements for this quest and draw a roadmap to obtain an evolution law satisfying all of them. Our nomenclature is shown in Tables 1 and 2.
@@
Several papers have claimed the successful reproduction of the three experimentally established behaviors mentioned above, but always within a limited range of variables; their complete reconciliation is impossible (§4.1). As mentioned above, the aging law and the slip law coincide around the steady states, so reconciling them is effortless around steady states. Nontrivial is the regime far from steady states, and a few studies succeeded in explaining far-from-steady-state behaviors. One outstanding proposal is the composite law (eq. 74) by Kato & Tullis (2001). However, the coincidences between those laws and experimental behaviors are all heuristic. Namely, they are not deduced from those experimental requirements, unlike the aging law and the slip law. Since the experimental requirement from VS and SS tests is exactly the slip law, which is besides supported by the stationary-loadpoint SHS, those heuristic laws cannot withstand the questions whether their predictions differing from the slip law are the case. Existing observations in VS and stationary-loadpoint SHS tests consistently support the slip law, even at slip rates almost the same as plate convergence rates (Bhattacharya et al., 2022). In other words, in the aging vs. slip problem context, the true importance lies in the constraint on reasonable deviations from the slip law: the constraint on how to joint frictional slip strengthening/weakening (the slip law) and the aging of stationary contacts (the aging law). Heuristics proposing many hows do not answer this question of determining actual how, regardless of the quality of fit.
@@
In §4.1, we prove that simultaneous satisfaction of the three canons is mathematically impossible. This incompatibility among the canons is the fundamental answer to why the aging and slip laws conflict. From this point forward, we pursue satisfying the three canons in the variable range where experiments have confirmed the canons and prove that this is possible if the aging law and the slip law are switched according to the value of the state relative to its steady-state value (eq. 86). Although we first introduce eq. (86) heuristically in §4.2.1 (eq. 84), we later rederive it deductively in §4.2.2 and again in §5.1 (eq. 125) with further generalization. When the state evolution law is premised to be a function of only the current slip rate and the current interface state, the evolution law satisfying eq. 125 is the unique choice to satisfy the three canons. A few examples of the concrete function form are presented in §4.3. Lastly, we extend the present results for plausible minor modifications of the canons (§4.4 and §4.5).
@@
Table 1: Nomenclature. The numbers indicate the equation number where the associated parameters and functions appear.
@@
$\Phi$ [N/m2] frictional strength; serves as a state variable (6–8)
@@
$L$ [m] $\theta/V$ at steady states (29)
@@
$V_{\rm c}$ [m/s] possible $V$ value s.t. $V\mathbin{\begin{smallmatrix}\gg\\ \ll\end{smallmatrix}}V_{\rm c}\Leftrightarrow f_{t}\mathbin{\begin{smallmatrix}\ll\\ \gg\end{smallmatrix}}f_{\delta}$ (74)
$\beta$ possible $\Omega$ value s.t. $\Omega\mathbin{\begin{smallmatrix}\gg\\ \ll\end{smallmatrix}}\beta\Leftrightarrow f_{t}\mathbin{\begin{smallmatrix}\ll\\ \gg\end{smallmatrix}}f_{\delta}$ (83)
@@ 2.1 Conventional view
Suppose a frictional interface between two elastic blocks under shear stress. This paper assumes mechanical equilibrium, neglecting inertial effects. Sliding or not, the stress continuity in the continuum host media proximate to the interface guarantees that the shear stress applied to the interface always equals frictional shear resistance. Thus, we use the same variable $\tau$ to express the magnitude of these two physical quantities. When the distinction is relevant in context, we refer to it as the applied shear stress or frictional resistance. Note, as customary in the field, we may refer to the frictional resistance as the frictional stress. The present paper concerns only the situations where the shear stress and slip rate directions are invariant; $\tau$ and $V$ denote their absolute values. Also, we assume normal stress $\sigma$ is constant.
@@
Among the existing evolution laws, the aging, slip, and composite (a hybrid of the aging and slip laws, Kato & Tullis, 2001) laws are deeply relevant to the present derivation of new evolution laws. We will explain their workings in light of the canonical behaviors in §3.2, 3.3, and 4.1.
Logically, there is no a priori reason to doubt the evolution law only: eq. (1) may be at fault as well (Nakatani, 2003; Baumberger & Caroli, 2006; Barbot, 2019). However, as previous authors did, we proceed by premising eq. (1). Although we add a minor modification of eq. (1) to deal with the lower limit of the frictional strength (§2.3, 4.5), it is irrelevant to our goal of simultaneous reproduction of the three canons.
@@ 2.2 Separation of the constitutive law and the state-evolution law
Thus, in the notation with $\Phi$, the RSF law separates into the rheological constitutive law (eq. 7, or equivalently 8) describing the instantaneous relationship between $\tau$ and $V$, conditioned on the parameter $\Phi$ (frictional strength of the interface at the moment), and the equation describing the time evolution of the frictional strength $\Phi$ (eq. 13). This set is mathematically equivalent to the conventional set of eqs. (1) and (3) through the variable conversion (eqs. 11 and 12). However, we need to be cautious of the fact that the state variable $\theta$ appears in eq. (1) in the form of $B\ln\theta$, where $B\ln(\cdot)$ is tied to some experimental facts regarding the evolution of $\Phi$. In consequence of using $\theta$-notation, the information on how the variation in the interface strength $B\ln(\theta/\theta_{*})$ occurs spans the two equations of eqs. (1) and (3). The constant $B$, appearing in eq. (1), is the quantity rooted in the two observations of log-V dependence of steady-state dynamic friction (§3.1) and log-t healing (§3.3), meaning that eq. (1) already premises specific knowledge of the time evolution of strength. Thus, eq. (1) in the $\theta$-notation is customarily bound to specific notions about the time evolution of strength. For us to pursue the phenomenological development of the strength evolution law without resorting to assumptions about the underlying physics, unsuited is the $\theta$-notation, which was introduced through the physical intuition that ascribes frictional strength variations to log-t healing and its erasure by slip (Dieterich, 1979). By contrast, in the rewritten RSF (the set of eqs. 7 and 13), where the frictional strength $\Phi$ serves as the state variable, the present subject of how the frictional interface evolves is solely taken care of by the evolution law (eq. 13) of $\Phi$.
Thus, our derivation (§3 and 4) of the evolution law shall start with describing the experimental requirements in the $\Phi$-notation, where the evolution law (eq. 13) well separates from the constitutive law (eq. 7). Naturally, the derived equations and findings are valid in any RSF formulations, including the $\theta$-notation (the set of eqs. 1 and 3) convertible via eqs. (11) and (12). We also show derived results in the $\theta$-notation to secure usability in this widely used notation.
@@ 2.3 Lower limit of frictional strength
When the interface is kept sliding at a fixed slip rate, frictional strength converges to a steady-state level $\Phi_{\rm SS}(V)$ specific to the slip rate. This convergence to a steady state has been well observed over various conditions in laboratories, thus premised in this study. This observation implies $\dot{\Phi}\gtrless 0$ when $\Phi\lessgtr\Phi_{\rm SS}(V)$, which in turn leads to $\dot{\Phi}>0$ whenever $\Phi$ is less than $\min_{\rm V}\Phi_{\rm SS}$, the lower limit of the steady-state strength. Therefore, we see that $\min_{\rm V}\Phi_{\rm SS}$ (denoted as $\Phi_{\rm X}$ hereafter) is also the lower limit of the frictional strength as long as we consider situations under fixed normal stress, as we declared at the beginning of §2.1. Namely,
@@
where $\theta_{\rm X}$ denotes the level of $\theta$ below which $\Phi$ shows little logarithmic dependence on $\theta$. According to the contact-age interpretation of $\theta$, $\theta_{\rm X}$ represents the short-time cutoff of the contact-age dependence of $\Phi$, and $\Phi_{\rm X}$ represents the minimum $\Phi$ value achieved at the short-time limit of the contact time $\theta\searrow 0$. Fig. 1 illustrates eq. (15); $\Phi$ loses the $\theta$ dependence for $\theta\ll\theta_{\rm X}$ and equals $\Phi_{\rm X}$ at $\theta=0$. Equation (15) introduces the strength minimum while retaining the definition range of $\theta$ ($0\leq\theta<\infty$) in the conventional evolution laws, such as the aging and slip laws.
@@ 3 Experimental requirements and empirical constraints on evolution laws
In §3.13.3, we formulate each canonical behavior of the frictional strength $\Phi$ in SS, VS, and SHS tests (Ruina, 1983; Dieterich, 1979) to derive constraints on evolution laws in mathematical forms, which we use in the present development of better evolution laws (§4). To focus on the role of the friction law, we consider ideal experiments where the loading apparatus is very stiff. Substantial consensus exists in the field on the canonical behaviors we formulate in this section. These behaviors have been the target of earlier works of the same purpose, i.e., finding an evolution law capable of reproducing all those canonical behaviors simultaneously; the present paper is trying to take part in this quest.
@@
We are aware that $\Phi$ can evolve also in response to the changes in normal stress $\sigma$ and shear stress $\tau$, but the present work does not consider them. The shear stress effect (Nagata et al., 2012) does not seem directly related to the difficulty of reproducing the evolutions in VS and SHS tests simultaneously, the main issue the present paper attempts to resolve. Although Nagata et al. (2012) pointed out that their evolution law, where the explicit effect of $\tau$ has been added to the aging law, approximately reproduced the symmetric evolutions following the positive and negative velocity steps, this success is limited to small-magnitude steps—Bhattacharya et al. (2015) have demonstrated that Nagata’s law, with the laboratory-derived magnitude of the shear stress effect (Nagata et al., 2012), predicts an asymmetric response if the velocity is stepped by as large as three orders of magnitude. Meanwhile, the normal stress effect (Linker & Dieterich, 1992) is beyond the scope of the present work. In any case, these stress effects reported so far are represented as additional stress-dependent evolution terms to be added to eq. (19). Hence, the presently developed evolution law, which has the form of eq. (19), can be extended to include those stress effects as additional terms.
@@
To start formulating canons, we introduce idealizations with respect to the experimental settings and their reasoning. The friction law describes a nature of the frictional interface, thus determined by the properties of the interface and free from the experimental apparatus. The RSF premises this natural expectation, known as the spatial locality of physical laws, and the RSF users correct the influence of machine stiffness and unwanted machine motions by using spring-slider models when fitting the RSF to experimental data. Conversely, since the friction law is independent of anything outside the interface, here we are allowed to use the idealized machine property and idealized machine motions when formulating laws. The derived law under such a premise must not depend on the apparatus, that is, designed to be applicable to actual cases with finite machine stiffness and unwanted machine motions. We follow those basic concepts of the RSF, and in formulating the canons, we suppose that ram motions are instantaneous, much faster than the time scale of interest, but within the quasistatic regime. In formulating the VS canon, we further assume the stiffness is sufficiently high. However, note that log-t healing during stationary-loadpoint SHS is affected by finite stiffness as $\tau(t)$ and $V(t)$ depend on $k$, while they are not affected by $k$ during constant-$\tau_{\rm hold}$.

### v3

@@ 3 Graduate School of Science
The aging law and the slip law are two representative evolution laws of the rate- and state-dependent friction (RSF) law, based on canonical behaviors in three types of laboratory experiments: slide-hold-slide (SHS), velocity-step (VS), and steady-state (SS) tests. The aging law explains the SHS canon but contradicts the VS canon, and vice versa for the slip law. The later proposed composite law, which switches these two laws according to the slip rate $V$, explains both canons but contradicts the SS canon. The present study constructs evolution laws satisfying all three canons throughout the range of variables where experiments have confirmed the canons. By recompiling the three canons, we have derived constraints on the evolution law and found that the evolution rates in the strengthening phases of the SHS and VS canons are so different that complete reconciliation throughout the entire range of variables is mathematically impossible. However, for the limited range of variables probed by experiments so far, we have found that the SHS and VS canons can be reconciled without violating the SS canon by switching the evolution function according to $\Omega$, the ratio of the state $\theta$ to its steady-state value $\theta_{\rm SS}$ for the instantaneous slip rate. We could generally show that, as long as the state evolution rate $\dot{\theta}$ depends only on the instantaneous values of $V$ and $\theta$, simultaneous reproduction of the three canons, throughout the experimentally confirmed range, requires the aging-law-like evolution for $\Omega$ sufficiently below a threshold $\beta$ and the slip-law-like evolution for $\Omega$ sufficiently above $\beta$. The validity of the canons in existing experiments suggests $\beta\lesssim 0.01$.
@@ 1 Introduction
Formulas that describe the consensual frictional behaviors in three representative types of experiments (eq. 26, eq. 32, and eqs. 48 and 49) have been pursued since Dieterich (1979). Dieterich’s formulation (1979) is primarily motivated by SS and SHS tests by Dieterich (1972) and Dieterich (1978). It was reorganized into a differential equation (eq. 4) by Ruina (1983), called the aging law (Dieterich law) today. However, the aging law contradicts VS tests. By linear expansion of the aging law around the steady states, Ruina (1983) proposed another differential equation (eq. 5) called the slip law (Ruina law) today. Ruina (1983) pointed out that the slip law is a straightforward translation of VS and SS behaviors into a differential equation. However, the slip law fails to describe the log-time healing observed under low shear stresses where slip is negligible, although it explains another, more popular type of SHS (the stationary-loadpoint SHS, §3.3.1) where shear stress remains pretty high, allowing for significant slip to continue in its quasi-static hold period. Thus, neither law has succeeded in compiling the different types of experiments. This long-standing discrepancy, the aging vs. slip problem (Beeler et al., 1994; Ampuero & Rubin, 2008), is the background of the present paper. Before starting the lengthy mathematical derivation, we summarize the requirements for this quest and draw a roadmap to obtain an evolution law satisfying all of them. Our nomenclature is shown in Tables 1 and 2.
@@
Several papers have claimed the successful reproduction of the three experimentally established behaviors mentioned above, but always within a limited range of variables; their complete reconciliation is impossible (§4.1). As mentioned above, the aging law and the slip law coincide around the steady states, so reconciling them is effortless around steady states. Nontrivial is the regime far from steady states, and a few studies succeeded in explaining far-from-steady-state behaviors. One outstanding proposal is the composite law (eq. 74) by Kato & Tullis (2001). However, the coincidences between those laws and experimental behaviors are all heuristic. Namely, they are not deduced from those experimental requirements, unlike the aging law and the slip law. Since the experimental requirement from VS and SS tests is exactly the slip law, which is besides supported by the stationary-loadpoint SHS, those heuristic laws do not withstand the questions whether their predictions differing from the slip law are the case. Existing observations in VS and stationary-loadpoint SHS tests consistently support the slip law, even at slip rates almost the same as plate convergence rates (Bhattacharya et al., 2022). In other words, in the aging vs. slip problem context, the true importance lies in the constraint on reasonable deviations from the slip law: the constraint on how to joint frictional slip strengthening/weakening (the slip law) and the aging of stationary contacts (the aging law). Heuristics proposing many hows do not answer this question of determining actual how, regardless of the quality of fit.
@@
In §4.1, we prove that simultaneous satisfaction of the three canons is mathematically impossible. This incompatibility among the canons is the fundamental answer to why the aging and slip laws conflict. From this point forward, we pursue satisfying the three canons in the variable range where experiments have confirmed the canons and prove that this is possible if the aging law and the slip law are switched according to the value of the state relative to its steady-state value (eq. 86). Although we first introduce eq. (86) heuristically in §4.2.1 (eq. 84), we later rederive it deductively in §4.2.2 and again in §5.1 (eq. 124) with further generalization. When the state evolution law is premised to be a function of only the current slip rate and the current interface state, the evolution law satisfying eq. (124) is the unique choice to satisfy the three canons. A few examples of the concrete function form are presented in §4.3. Lastly, we extend the present results for plausible minor modifications of the canons (§4.4 and §4.5).
@@
Table 1: Nomenclature. The numbers indicate the equations where the associated parameters and functions appear.
@@
$\Phi$ [N/m2] frictional strength; serves as a state variable (68)
@@
$L$ [m] $V\theta$ at steady states (28)
@@
$V_{\rm c}$ [m/s] possible $V$ value s.t. $V\mathbin{\begin{smallmatrix}\gg\\ \ll\end{smallmatrix}}V_{\rm c}\Leftrightarrow f_{t}\mathbin{\begin{smallmatrix}\ll\\ \gg\end{smallmatrix}}f_{\delta}V$ (74)
$\beta$ possible $\Omega$ value s.t. $\Omega\mathbin{\begin{smallmatrix}\gg\\ \ll\end{smallmatrix}}\beta\Leftrightarrow f_{t}\mathbin{\begin{smallmatrix}\ll\\ \gg\end{smallmatrix}}f_{\delta}V$ (83)
@@ 2.1 Conventional view
Suppose a frictional interface between two elastic blocks under shear stress. The present study assumes mechanical equilibrium, neglecting inertial effects. Sliding or not, the stress continuity in the continuum host media proximate to the interface guarantees that the shear stress applied to the interface always equals frictional shear resistance. Thus, we use the same variable $\tau$ to express the magnitude of these two physical quantities. When the distinction is relevant in context, we refer to it as the applied shear stress or frictional resistance. Note, as customary in the field, we may refer to the frictional resistance as the frictional stress. The present paper concerns only the situations where the shear stress and slip rate directions are invariant; $\tau$ and $V$ denote their absolute values. Also, we assume normal stress $\sigma$ is constant.
@@
Among the existing evolution laws, the aging, slip, and composite (a hybrid of the aging and slip laws, Kato & Tullis, 2001) laws are deeply relevant to the present derivation of new evolution laws. We will explain their workings in light of the canonical behaviors in §3.2, §3.3, and §4.1.
Logically, there is no a priori reason to doubt the evolution law only: eq. (1) may be at fault as well (Nakatani, 2003; Baumberger & Caroli, 2006; Barbot, 2019). However, as previous authors did, we proceed by premising eq. (1). Although we add a minor modification of eq. (1) to deal with the lower limit of the frictional strength (§2.3 and §4.5), it is irrelevant to our goal of simultaneous reproduction of the three canons.
@@ 2.2 Separation of the constitutive law and the state-evolution law
Thus, in the notation with $\Phi$, the RSF law separates into the rheological constitutive law (eq. 7, or equivalently, 8) describing the instantaneous relationship between $\tau$ and $V$, conditioned on the parameter $\Phi$ (frictional strength of the interface at the moment), and the equation describing the time evolution of the frictional strength $\Phi$ (eq. 13). This set is mathematically equivalent to the conventional set of eqs. (1) and (3) through the variable conversion (eqs. 11 and 12). However, we need to be cautious of the fact that the state variable $\theta$ appears in eq. (1) in the form of $B\ln\theta$, where $B\ln(\cdot)$ is tied to some experimental facts regarding the evolution of $\Phi$. In consequence of using $\theta$-notation, the information on how the variation in the interface strength $B\ln(\theta/\theta_{*})$ occurs spans the two equations of eqs. (1) and (3). The constant $B$, appearing in eq. (1), is the quantity rooted in the two observations of log-V dependence of steady-state dynamic friction (§3.1) and log-t healing (§3.3), meaning that eq. (1) already premises specific knowledge of the time evolution of strength. Thus, eq. (1) in the $\theta$-notation is customarily bound to specific notions about the time evolution of strength. For us to pursue the phenomenological development of the strength evolution law without resorting to assumptions about the underlying physics, unsuited is the $\theta$-notation, which was introduced through the physical intuition that ascribes frictional strength variations to log-t healing and its erasure by slip (Dieterich, 1979). By contrast, in the rewritten RSF (the set of eqs. 7 and 13), where the frictional strength $\Phi$ serves as the state variable, the present subject of how the frictional interface evolves is solely taken care of by the evolution law (eq. 13) of $\Phi$.
Thus, our derivation (§3 and §4) of the evolution law shall start with describing the experimental requirements in the $\Phi$-notation, where the evolution law (eq. 13) well separates from the constitutive law (eq. 7). Naturally, the derived equations and findings are valid in any RSF formulations, including the $\theta$-notation (the set of eqs. 1 and 3) convertible via eqs. (11) and (12). We also show derived results in the $\theta$-notation to secure usability in this widely used notation.
@@ 2.3 Lower limit of frictional strength
When the interface is kept sliding at a fixed slip rate, frictional strength converges to a steady-state level $\Phi_{\rm SS}(V)$ specific to the slip rate. This convergence to a steady state has been well observed over various conditions in laboratories, thus premised in the present study. This observation implies $\dot{\Phi}\gtrless 0$ when $\Phi\lessgtr\Phi_{\rm SS}(V)$, which in turn leads to $\dot{\Phi}>0$ whenever $\Phi$ is less than $\min_{\rm V}\Phi_{\rm SS}$, the lower limit of the steady-state strength. Therefore, we see that $\min_{\rm V}\Phi_{\rm SS}$ (denoted as $\Phi_{\rm X}$ hereafter) is also the lower limit of the frictional strength as long as we consider situations under fixed normal stress, as we declared at the beginning of §2.1. Namely,
@@
where $\theta_{\rm X}$ denotes the level of $\theta$ below which $\Phi$ shows little logarithmic dependence on $\theta$. According to the contact-age interpretation of $\theta$, $\theta_{\rm X}$ represents the short-time cutoff of the contact-age dependence of $\Phi$, and $\Phi_{\rm X}$ represents the minimum $\Phi$ value achieved at the short-time limit of the contact time $\theta\searrow 0$. Figure 1 illustrates eq. (15); $\Phi$ loses the $\theta$ dependence for $\theta\ll\theta_{\rm X}$ and equals $\Phi_{\rm X}$ at $\theta=0$. Equation (15) introduces the strength minimum while retaining the definition range of $\theta$ ($0\leq\theta<\infty$) in the conventional evolution laws, such as the aging and slip laws.
@@ 3 Experimental requirements and empirical constraints on evolution laws
In §3.13.3, we formulate each canonical behavior of the frictional strength $\Phi$ in SS, VS, and SHS tests (Ruina, 1983; Dieterich, 1979) to derive constraints on evolution laws as empirical formulae, which we use in the present development of better evolution laws (§4). Substantial consensus exists in the field on the canonical behaviors we formulate in this section. These behaviors have been the target of earlier works of the same purpose, i.e., finding an evolution law capable of reproducing all those canonical behaviors simultaneously; the present paper is trying to take part in this quest.
@@
We are aware that $\Phi$ can evolve also in response to the changes in normal stress $\sigma$ and shear stress $\tau$, but the present work does not consider them. The shear stress effect (Nagata et al., 2012) does not seem directly related to the difficulty of reproducing the evolutions in VS and SHS tests simultaneously, the main issue the present paper attempts to resolve. Although Nagata et al. (2012) noted that their evolution law, where the explicit effect of $\tau$ has been added to the aging law, approximately reproduced the symmetric evolutions following the positive and negative velocity steps, this success is limited to small-magnitude steps—Bhattacharya et al. (2015) have demonstrated that the Nagata law, with the laboratory-derived magnitude of the shear stress effect (Nagata et al., 2012), predicts an asymmetric response if the velocity is stepped by as large as three orders of magnitude. Meanwhile, the normal stress effect (Linker & Dieterich, 1992) is beyond the scope of the present work. In any case, these stress effects reported so far are represented as additional stress-dependent evolution terms to be added to eq. (19). Hence, the presently developed evolution law, which has the form of eq. (19), can be extended to include those stress effects as additional terms.
@@
To start formulating canons, we introduce idealizations with respect to the experimental settings and their reasoning. The friction law describes a nature of the frictional interface, thus determined by the properties of the interface and free from the experimental apparatus. The RSF premises this natural expectation, known as the spatial locality of physical laws, and the RSF users accounts for the influence of machine stiffness and unwanted machine motions by using spring-slider models when fitting the RSF to experimental data. Conversely, since the friction law itselef is independent of anything outside the interface, here we are allowed to use the idealized machine property and idealized machine motions when formulating laws, without loss of applicability to actual experiments. We follow these basic concepts of the RSF, and in formulating the canons, we suppose that ram velocity and force changes can be stepwise, much faster than the time scale of interest, but within the quasistatic regime. In formulating the VS canon, we further assume the stiffness is sufficiently high. Meanwhile, note that log-t healing during stationary-loadpoint SHS is affected by finite stiffness $k$ as $\tau(t)$ and $V(t)$ depend on $k$, while they are not affected by $k$ during constant-$\tau$ holds (§3.3.1).